OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

hartnn (hartnn):

add 4 on both sides first

OpenStudy (anonymous):

Okay 2x + 8 = 20 ?

hartnn (hartnn):

not like that |2x+6| -4+4 = 20+4 |2x+6| = 24 got that ?

OpenStudy (anonymous):

Oh okay! that makes more sense sooo, 2x +6 = 24 ?

hartnn (hartnn):

since we have absolute sign, we would have both 2x+6 = 24 or 2x+6 = -24 can you solve both of these to get 2 values of x ?? subtract 6 from both sides,....

OpenStudy (anonymous):

I don't get it....

hartnn (hartnn):

2x+6 =24 can you find x from here ?

OpenStudy (anonymous):

i mean 2x = 24

hartnn (hartnn):

2x+6 = 24 2x = 24-6 = 18 2x=18 x=9 :) now find x from 2x+6 = -24

OpenStudy (anonymous):

Well I put -15 times two plus six and got -24 ? does that work or is there no solution ?

hartnn (hartnn):

2x+6 = -24 2x = -30 x=-15 works! so its, 9,-15

OpenStudy (anonymous):

yay its starting to make sense!

hartnn (hartnn):

glad to hear that! :)